OpenStudy (anonymous):

What is the factor of 3t^2 - 10t?

OpenStudy (haseeb96):

t(3t-10) is the factors of it. and this is correct

OpenStudy (anonymous):

How about when I'm going to find t in 0 = 3t^2 - 10t?

OpenStudy (haseeb96):

just t(3t-10)=0 either t=0 or 3t-10=0 t=0 or 3t=10 t=0 or t=10/3 this is the correct answer and if u want to check that either yur answer is correct nor not . then put first t=0 in 3t^2-10t=0 then u will find 0-0=0 0=0 and when the both sides values will be equal then yur answer is correct

OpenStudy (haseeb96):

have u got it @mtalampas21

OpenStudy (anonymous):

Yah, thanks :)